OpenStudy (anonymous):

m=((Y-Y_0))/((X-X_0)) and i need to solve for x

OpenStudy (anonymous):

like this? \[\huge m=\frac{y-y_0}{x-x_0} \] solve for x...?

OpenStudy (anonymous):

yes thats exactly it! i just didnt know how to insert the subscripts!

OpenStudy (anonymous):

multiply both sides by \(\large x-x_0 \) to get: \[\large m(x-x_0)=y-y_0 \] then divide both sides by m to get \[\large x-x_0=\frac{y-y_0}{m} \] then finally add \(\large x_0 \) to both sides: \[\large x=\frac{y-y_0}{m}+x_0 \]

OpenStudy (anonymous):

how would i verify that result on a test?

OpenStudy (anonymous):

whadda ya meen?

OpenStudy (anonymous):

the test asks that i verify my result and i have no idea how i would verify my answer. what you put is exactly how i answered it but i dont understand how to verify the problem

OpenStudy (anonymous):

this is one way form the what is called point-slope form.... when you mean verify, i think that maybe you will be given a point and it will satisfy that equation we ended up with... ???

OpenStudy (anonymous):

okay and i have another question and i was hoping you could walk me thru.. i need to solve (1/A)-(1/B)=(1/C) and i need to solve for A

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

dammit.... i got an "aw snap" message...

OpenStudy (anonymous):